# RadioLLM: Introducing Large Language Model into Cognitive Radio via Hybrid Prompt and Token Reprogrammings

### v1

@@ C.3.3 Structural Similarity Index (SSIM)
Appendix D Model Analysis
D.1 Q&A Analysis on Applying LLMs to Radio Signal Feature Extraction
The application of large language models (LLMs) to radio signal feature extraction is an emerging research direction that combines advancements in natural language processing (NLP) with signal processing techniques. Below is a detailed analysis of the challenges, advantages, and considerations involved in this novel approach.
1. Why Use LLMs for Radio Signal Feature Extraction?
LLMs, such as GPT-2, BERT, and LLaMA, are designed to handle sequence data, making them versatile for tasks beyond textual data. Radio signals can also be represented as sequences (e.g., time-domain waveforms or frequency-domain representations). LLMs can be adapted to radio signal feature extraction due to their ability to capture long-range dependencies and complex patterns, which are crucial for analyzing radio signals.
Advantages:
  1. 1.
    Sequence Modeling: LLMs excel at processing sequential data, which aligns well with the inherent sequential nature of radio signals.
  2. 2.
    Contextual Understanding: LLMs can learn contextual relationships in signal features, such as frequency bands, modulation types, and time-domain patterns.
  3. 3.
    Transfer Learning: Pretrained LLMs on large datasets (e.g., self-supervised learning) can be fine-tuned for radio signal tasks with limited labeled data, reducing the need for large-scale datasets.
  4. 4.
    Robustness to Noise: LLMs can learn to generalize across noisy data, making them effective for real-world radio environments, where interference and noise are common.
Example Applications:
  1. 1.
    Modulation Recognition: Automatically identifying modulation schemes (e.g., AM, FM, QPSK) from signal data.
  2. 2.
    Spectrum Analysis: Extracting and classifying spectral features for interference detection or signal identification.
  3. 3.
    Signal Detection: Identifying weak signals buried in noise or interference using learned patterns.
2. Key Challenges in Applying LLMs to Radio Signals
Applying large language models (LLMs) to radio signal processing presents several unique challenges stemming from the fundamental differences between radio signals and text data. One of the primary difficulties lies in representing radio signals in a format compatible with LLMs. Common representations include raw IQ (In-phase and Quadrature) samples, which are complex-valued time-domain data; spectrograms, which transform signals into time-frequency representations; and feature vectors, which capture extracted characteristics like spectral peaks or bandwidth. Selecting the most suitable representation is critical, as it directly impacts how well the LLM can learn and process the signal’s key features.
Another challenge is the high dimensionality of radio signals compared to text. High-bandwidth signals or long observation windows can result in sequences that are far longer than those typically encountered in NLP tasks. Feeding such long sequences into LLMs can significantly increase computational costs and memory requirements. Additionally, reducing the sequence length through downsampling or truncation risks the loss of essential information, which may degrade performance on tasks such as modulation recognition or spectrum analysis. This trade-off between computational efficiency and information retention is a key limitation when applying LLMs directly to raw signal data.
Finally, radio signals possess unique features that LLMs are not inherently designed to handle. For example, phase and amplitude relationships are critical for understanding modulation schemes but are not explicitly modeled in standard LLM architectures. Real-world radio environments also introduce noise and interference, which can act as adversarial inputs, further complicating the extraction of meaningful features. Moreover, the scarcity of large-scale, labeled radio signal datasets limits the ability to train or fine-tune LLMs effectively, unlike in NLP, where massive text corpora are readily available. These challenges highlight the need for specialized adaptations in data representation, model design, and training strategies to fully leverage the potential of LLMs for radio signal feature extraction.
D.1.1 LLM goes beyond dialogue: the potential of signal feature extraction
LLMs have traditionally been associated with natural language processing tasks, such as dialogue generation and text understanding. However, their versatility extends far beyond these applications. Recent advancements demonstrate that LLMs can be effectively applied to domains outside of text, including signal feature extraction, where they leverage their contextual understanding capabilities to identify patterns and relationships in complex datasets.
In the context of the RadioLLM framework, LLMs can process signal-related data by treating it as a sequence of features analogous to text tokens. This perspective enables models such as GPT-2, BERT, and LLaMA3 to extract meaningful features from signals, classify them, and even infer signal properties that are challenging to detect using traditional machine learning methods. For example, by training on domain-specific embeddings, LLMs can identify nuanced relationships between signal amplitude, frequency components, and modulation patterns, which are critical for tasks like signal classification or interference detection.
The potential of LLMs in signal processing is further enhanced by their scalability and ability to learn from large datasets. Unlike conventional methods, which often rely on handcrafted features, LLMs can automatically learn hierarchical representations, reducing the need for manual feature engineering. This makes them particularly well-suited for dynamic environments, where signals can vary significantly over time or across contexts. However, challenges such as high computational costs, long inference times, and the need for domain-specific fine-tuning remain barriers to widespread adoption.
Future research could explore how lightweight LLM architectures, such as distilled or quantized models, can be optimized for real-time signal processing tasks without sacrificing accuracy. Additionally, hybrid approaches that combine LLMs with traditional signal processing techniques could unlock new possibilities for efficient and robust feature extraction. This illustrates that the utility of LLMs extends far beyond dialogue systems, paving the way for innovative applications in signal analysis and beyond.
D.1.2 Advancing CRT: How Radiollm Differs from Prior Work
RadioLLM represents a significant leap in cognitive radio technology (CRT), standing apart from prior approaches through its innovative design and comprehensive capabilities. Unlike traditional models that typically focus on a single task or rely on limited datasets, RadioLLM employs a hybrid and diverse dataset for pre-training, enabling it to capture a broader spectrum of signal features. This hybrid dataset incorporates a variety of signal types and conditions, allowing the model to generalize effectively across multiple scenarios.
Moreover, RadioLLM distinguishes itself by addressing multiple downstream tasks—classification, imputation, and denoising—within a unified framework. This all-in-one capability reduces the need for task-specific models, streamlining the signal processing pipeline while maintaining high performance across tasks. By leveraging the powerful feature extraction capabilities of pre-trained large language models, RadioLLM bridges the gap between linguistic intelligence and radio signal processing, offering a versatile solution for complex signal environments.
Through this novel approach, RadioLLM not only advances CRT but also establishes a new benchmark for multi-task adaptability and efficiency in the field.

### v3

@@ IV-E3 SSIM
IV-F Comparison with RSC methods
TABLE I: Comparison of Methods on Classification Tasks. BOLD indicates the best performance, and UNDERLINED indicates the second-best performance.
Methods RML16A RML16B RML16C RML22 RML18A Wi-Fi ADS-B
OA(%) Kappa OA(%) Kappa OA(%) Kappa OA(%) Kappa OA(%) Kappa OA(%) Kappa OA(%) Kappa
HCGDNN [31] 54.30 0.4973 55.30 0.5033 65.02 0.6097 56.93 0.5215 31.29 0.2831 15.67 0.1004 44.66 0.4434
PETCGDNN [32] 51.09 0.4620 50.91 0.4546 62.12 0.5771 56.05 0.5116 28.70 0.2560 34.59 0.3023 97.47 0.9746
MCLDNN [33] 52.84 0.4812 52.45 0.4716 65.63 0.6130 53.28 0.4809 42.05 0.3953 19.41 0.1403 65.12 0.6493
CVCNN [34] 54.15 0.4957 54.06 0.4896 65.49 0.6142 54.57 0.4953 34.79 0.3196 11.44 0.0553 97.49 0.9747
SFS-SEI [35] 54.13 0.4954 54.17 0.4907 65.45 0.6143 53.49 0.4832 36.96 0.3422 11.45 0.0555 97.50 0.9749
ICAMCNET [36] 53.55 0.4891 53.43 0.4825 63.17 0.5892 54.23 0.4914 36.50 0.3374 11.34 0.0543 99.09 0.9908
CVSRN [37] 47.56 0.4232 51.50 0.4611 64.41 0.6021 50.09 0.4454 35.74 0.3295 11.72 0.0583 94.55 0.9452
TcssAMR [38] 56.53 0.5218 55.98 0.5109 66.21 0.6224 58.17 0.5353 47.35 0.4506 32.86 0.2873 61.89 0.6177
SemiAMC [39] 55.32 0.5076 56.17 0.5129 64.62 0.6056 58.08 0.5342 47.66 0.4539 29.36 0.2502 70.42 0.7034
RadioLLM 58.10 0.5391 58.35 0.5372 68.19 0.6441 59.39 0.5488 52.03 0.4994 35.41 0.3110 90.58 0.9054
We conducted extensive performance evaluations of our proposed model, RadioLLM, against several representative RSC approaches, including SL methods such as HCGDNN [31], PETCGDNN [32], MCLDNN [33], ICMACNET [36], CVCNN [34], SFS-SEI [35], CVSRN [37], and SSL methods such as SemiAMC [39] and TCSSAMR [38]. For methods without publicly available implementations, we carefully reconstructed network architectures based on original descriptions and retrained them using identical experimental conditions for fair comparisons. All experiments utilized multiple well-established datasets, specifically RML16A, RML16B, RML16C, RML22, RML18A, Wi-Fi, and ADS-B, under the 100-shot scenario, except for ADS-B where 10% of the data was used due to its limited size.
Table I summarizes the comprehensive performance metrics, including OA and Kappa coefficients for each method across all evaluated datasets under the 100-shot scenario. Clearly, RadioLLM demonstrates state-of-the-art performance on most datasets. Specifically, RadioLLM achieves the highest accuracy of 58.10%, 58.35%, 68.19%, 59.39%, 52.03%, and 35.41% on RML16A, RML16B, RML16C, RML22, RML18A, and Wi-Fi datasets, respectively. This substantial improvement highlights its robust ability to effectively leverage pre-trained knowledge for optimized feature extraction and transfer learning across different signal types.
Figure 4: Performance evaluation of RadioLLM across multiple datasets and SNR levels. Each column corresponds to one dataset: RML2016a, RML16B, RML16C, RML22, and RML18A. (a)–(e) show the OA of RadioLLM compared to other models under varying few-shot sample settings. (f)–(j) depict the OA under different SNR conditions with the few-shot number fixed at 100. (k)–(o) present the confusion matrices of RadioLLM on the corresponding datasets, illustrating detailed classification results across modulation classes.
On the RML16A dataset, the proposed model outperformed TcssAMR under the 100-shot setting, underscoring the benefits of pre-training for effective feature extraction. However, as illustrated in Fig. 4 (k), the model encounters difficulties in distinguishing between similar modulation types, such as 16-QAM and 64-QAM, as well as noise-sensitive signals like WBFM and AM-DSB. Notably, a substantial proportion of samples from several classes are misclassified as AM-SSB, with QPSK exhibiting a 26.4% misclassification rate into this category.
Furthermore, Fig. 4 (f) demonstrates that RadioLLM achieves superior performance at high SNR levels (greater than 0 dB), surpassing all baseline models. Even at low SNRs, it maintains competitive accuracy, albeit with performance slightly lower than MCLDNN under certain conditions. Analysis of the few-shot learning scenario, as shown in Fig. 4(a), reveals that the accuracy of all models increases with the number of training samples. Among them, MCLDNN displays the most pronounced improvement as the sample size grows. Although RadioLLM achieves comparable performance to traditional models in some extreme low-sample regimes, its advantage becomes increasingly apparent as more samples are provided, ultimately demonstrating outstanding performance as the training set expands.
On the RML16B dataset, the proposed RadioLLM model maintains outstanding performance, surpassing the runner-up SemiAMC by 2.18% in terms of OA, as shown in Table I. This result highlights the strong classification capability and the robust feature extraction capacity of the large-scale model, which benefits from both pre-training and architectural innovations. Moreover, RadioLLM achieves the highest Kappa coefficient, further indicating its superior consistency and reliability in multi-class signal classification tasks.
In terms of few-shot learning, Fig. 4 (b) reveals that MCLDNN is less adaptive to limited training data and thus exhibits relatively poor performance in low-sample scenarios. However, its accuracy increases sharply as the number of training samples grows. In contrast, RadioLLM, TcssAMR, and SemiAMC exhibit only marginal differences when trained with small sample sizes, with RadioLLM holding a slight advantage. This phenomenon can be attributed to the large parameter count of RadioLLM, which makes it challenging to optimize under extremely limited data. As the sample size increases, the performance gap widens, and RadioLLM’s advantage becomes more prominent, demonstrating its scalability and capacity to leverage larger datasets effectively.
Fig. 4 (g) further demonstrates that RadioLLM excels in high SNR conditions, outperforming other models by approximately 3%. At low SNRs, its performance remains competitive and comparable to baseline methods, indicating stable robustness under noisy environments. The confusion matrix in Fig. 4 (l) reflects patterns similar to those observed on RML16A, with challenges remaining in distinguishing between closely related modulation schemes such as 16-QAM and 64-QAM, as well as noise-sensitive signals like WBFM and AM-DSB. These findings suggest that, while RadioLLM is highly effective and generally robust, future work could focus on further enhancing its discrimination ability for highly similar or ambiguous classes.
On the RML16C dataset, the proposed RadioLLM model also achieves the best performance, outperforming all competing methods with an overall accuracy of 68.19% and a Kappa coefficient of 0.6441, as shown in Table I. In the few-shot learning setting, Fig. 4 (c) shows that all models benefit from increased sample size, with RadioLLM consistently maintaining a leading position as the number of training samples grows. As illustrated in Fig. 4 (h), RadioLLM exhibits robust performance across a wide range of SNRs, surpassing other models particularly at higher SNRs, which confirms its resilience to noise and superior classification capability under challenging conditions. Furthermore, the confusion matrix in Fig. 4 (m) demonstrates that RadioLLM achieves excellent class discrimination overall, with no severe misclassification except for notable confusion between AM-DSB and WBFM.
To further evaluate the generalization capability of the proposed model, we extended our experiments to the RML22 dataset, which shares similar characteristics with RML16A. Notably, we did not perform any pre-training on RML22, thereby providing a rigorous test of the model’s domain transferability. As shown in Table I, RadioLLM still achieves competitive performance, surpassing TcssAMR by 1.22% in OA and achieving the highest Kappa coefficient of 0.5488 among all compared methods. Regarding few-shot learning, Fig. 4 (d) demonstrates that RadioLLM steadily outperforms other models as the number of training samples increases, maintaining a consistent advantage across different sample size regimes. This result highlights the model’s strong adaptation ability and robustness, even when trained with limited data.
As depicted in Fig. 4 (i), RadioLLM maintains robust performance across a wide range of SNR levels, and clearly outperforms the baselines at higher SNRs, further confirming its resilience to noise and its capacity for generalization in unseen domains. The confusion matrix in Fig. 4 (n) shows that while RadioLLM achieves high accuracy for most classes, the main sources of misclassification are between QPSK and 8PSK, as well as between QAM16 and QAM64.
RadioLLM also outperformed competing models on more complex datasets. On RML18A, it managed increased task complexity better than MCLDNN, while on the Wi-Fi dataset, it achieved the highest performance among all methods tested, effectively handling intricate signal patterns. As shown in Fig. 4 (o), for the RML18A dataset, misclassifications among similar modulation types, such as 8PSK and BPSK, were infrequent. This highlights the effectiveness of transfer learning in distinguishing between closely related signal types. On the ADS-B dataset, while RadioLLM did not achieve the best performance, its results were still acceptable.
We further examine RadioLLM’s adaptability under varying few-shot scenarios, as depicted in Fig. 4 (a)-(e). RadioLLM consistently outperforms baseline methods across various few-shot settings, particularly evident when training samples are extremely limited (e.g., 1-shot and 5-shot). For instance, on the RML16A and RML16B datasets, RadioLLM maintains a significant accuracy advantage even with fewer than 10 labeled samples, underscoring its superior generalization capabilities. This advantage becomes more pronounced with increasing sample sizes, reflecting the effectiveness of the pre-trained feature representations obtained from large-scale data.
On more challenging datasets such as RML18A, which contains a large number of modulation classes, RadioLLM achieves a substantial performance improvement, outperforming the second-best method by 4.37% in OA and achieving the highest Kappa coefficient, as shown in Table I. In the few-shot regime, as depicted in Fig. 4 (e), RadioLLM maintains a consistent advantage of approximately 2% over the second-best model across varying sample sizes, demonstrating robust adaptation even with limited training data. Furthermore, as illustrated in Fig. 4 (j), RadioLLM exhibits significant gains in high SNR conditions, achieving around 10% higher accuracy than the next best method at SNR levels above 12 dB, highlighting its strong noise resilience.
The confusion matrix in Fig. 4 (o) reveals that the primary sources of misclassification occur between 4ASK and 8ASK, 16PSK and 32PSK, as well as among the higher-order QAM classes (64QAM, 128QAM, and 256QAM). Fig. 4 (f)-(j) details the performance of RadioLLM under varying SNRs. RadioLLM consistently exhibits superior robustness to noise compared with competing methods, especially in low-SNR scenarios. Specifically, RadioLLM significantly surpasses other methods in the challenging negative SNR range (below 0 dB), where accurate modulation recognition is inherently difficult. This robustness is particularly evident in datasets such as RML16C and RML22, highlighting RadioLLM’s capability to effectively extract meaningful signal features despite noisy conditions.
RadioLLM’s superior performance can be attributed to its effective integration of large-scale pre-training and fine-tuning strategy. By leveraging comprehensive representations learned from extensive signal data, RadioLLM efficiently adapts to diverse modulation classification tasks, exhibiting strong transfer learning capabilities. Moreover, its sophisticated architecture, which integrates layered feature extraction mechanisms with attention-based modeling, provides meaningful hierarchical representations that effectively discriminate among subtle differences in complex modulation schemes.
In summary, our extensive evaluations clearly demonstrate RadioLLM’s exceptional classification accuracy, robust generalization across diverse datasets, and resilience under challenging noise conditions. While certain confusion remains among closely related modulation types in highly noisy contexts, these results establish RadioLLM as a highly promising method for realistic radio signal classification tasks, offering significant advantages over contemporary approaches.
IV-G Performance Analysis of Baselines with Multi-Domain Pre-training
To further explore the benefits of large-scale and multi-domain learning, we conducted additional experiments in which baseline methods TcssAMR and SemiAMC were subjected to pre-training on multiple datasets. Specifically, both models were first pre-trained on an aggregated dataset composed of RML16A, RML16B, and RML16C, and subsequently fine-tuned individually on each of these datasets. This experimental design aims to assess whether exposure to diverse domain distributions during pre-training can enhance the models’ generalization capacity and downstream performance.
The experimental results are visualized in Fig. 5. As observed, multi-dataset pre-training generally leads to improvements across several evaluation metrics, such as OA and Kappa coefficient, compared to models trained from scratch on a single dataset. Both TcssAMR and SemiAMC exhibit higher performance on some datasets after multi-domain pre-training, indicating better feature transfer and increased robustness to domain shifts.
Figure 5: Comparison of OA for different methods on RML16A, RML16B, and RML16C. “All” denotes variants pre-trained jointly on all three datasets. The results illustrate the effect of multi-domain pre-training on model adaptability and generalization.
However, it is noteworthy that while multi-domain pre-training promotes performance gains in many scenarios, it also results in slight performance drops on certain datasets. This phenomenon is anticipated and can be attributed to the intrinsic characteristics of both the datasets and the models. The aggregated pre-training exposes the model to a broader range of data distributions, which enhances its ability to generalize across domains. Nevertheless, the same diversity may introduce domain-specific biases that are not fully aligned with the target dataset, leading to less optimal adaptation in some cases.
This result highlights a fundamental property of large-scale, pre-trained models: their adaptability to multiple datasets stems from their capacity to learn generalizable and transferable representations. Yet, the presence of both improvements and declines underscores the inherent trade-off between broad generalization and domain-specific optimization. The observed variance in performance reflects the model’s bias induced by pre-training, which can either align well with or deviate from the target domain characteristics.
In summary, our findings demonstrate that large-scale, multi-domain pre-training equips baseline models with enhanced adaptability and the ability to transfer knowledge across various scenarios. Nevertheless, achieving optimal performance in real-world signal processing tasks often requires further architectural innovations or domain-specific fine-tuning to fully leverage the strengths of foundation models while mitigating the risks of domain mismatch. This underscores the importance of both large-scale pre-training for broad applicability and targeted model design for robust signal classification in complex environments.
Figure 6: Comparison of denoising results on randomly selected high-SNR samples from the RML16A, RML16B, and RML16C datasets. For each dataset, the first column shows the original clean signal (Original), and the second column presents the corresponding signal after additive noise (Noise). The subsequent columns display the denoised results obtained by the SGFilter, DNCNet, and the proposed RadioLLM, respectively. RadioLLM demonstrates superior denoising performance, more effectively preserving the structure of the original signal under various conditions.
IV-H Comparison with Denoise methods
We comprehensively evaluated the proposed RadioLLM framework against both traditional signal processing and deep learning-based denoising models across three benchmark datasets: RML16A, RML16B, and RML16C. For methods lacking publicly available implementations, we re-implemented and retrained them in strict accordance with the protocols described in their original publications. During testing, to rigorously assess model robustness under realistic noise conditions, we injected additive noise into the test data, varying the SNR from 0dB to 10dB. This evaluation strategy ensures that all models are compared fairly under challenging and diverse noise environments, further highlighting the effectiveness of the proposed approach.
As illustrated in Fig. 6, a qualitative comparison is presented using randomly selected high-SNR samples from each dataset. The original clean signals (Original) are shown alongside their noisy counterparts (Noise), followed by the denoised results produced by SGFilter [40], DNCNet [41], and the proposed RadioLLM. Visually, RadioLLM is able to recover signal waveforms that are much closer to the original, effectively suppressing noise while preserving both the I and Q components’ structures. In contrast, both SGFilter and DNCNet either oversmooth the signal or fail to adequately eliminate noise, resulting in significant distortion or residual artifacts.
TABLE II: SSIM Values for Different Models Across Datasets.
Method RML16A RML16B RML16C
SGFilter [40] 0.782 0.821 0.777
DNCNet [41] 0.721 0.742 0.698
RadioLLM 0.838 0.893 0.846
Quantitative results, summarized in Table II, further validate these observations. Among all evaluated methods, RadioLLM consistently achieves the highest SSIM values across all datasets (0.838, 0.893, and 0.846 for RML16A, RML16B, and RML16C, respectively), indicating superior structural preservation and fidelity in the denoised outputs. SGFilter, despite being a classical smoothing technique, achieves the second-best SSIM scores, but its fixed and non-adaptive nature limits its effectiveness in handling the complex and diverse noise patterns typical in modern wireless signals. While SGFilter performs reasonably well on simpler noise profiles, it lacks the learning capability required for more sophisticated denoising tasks.
DNCNet, a deep learning-based model, exhibits lower SSIM scores across all datasets compared to both SGFilter and RadioLLM. This suggests that, although DNCNet benefits from neural network-based modeling, it struggles to fully capture the multi-scale temporal and frequency dependencies essential for robust denoising.
In contrast, RadioLLM leverages hierarchical relational modeling and frequency-aware fusion mechanisms, enabling more effective extraction of relevant signal features and superior generalization across diverse and challenging conditions. These results, corroborated by both qualitative and quantitative analyses, highlight the robustness and adaptability of RadioLLM, firmly establishing its superiority over both traditional and deep learning-based denoising baselines.
IV-I Ablation Studies
To thoroughly assess the effectiveness of the proposed HTRP and FAF modules, a series of controlled ablation experiments were conducted. These experiments systematically examined the individual and combined contributions of each module to the overall model performance, providing deeper insights into their respective roles. Unless explicitly stated otherwise, all ablation studies were performed on the RML16A dataset, following the experimental setup outlined in Section IV-C. The results are summarized in Table III.
TABLE III: Ablation Study Results for the HTRP and FAF Modules.
HTRP FAF OA (%) Kappa SSIM Inference Time(ms)
55.39 0.5097 0.805 1131.0
57.08 0.5283 0.833 766.5
57.25 0.5306 0.857 1149.4
58.10 0.5391 0.838 783.8
As presented in Table III, the baseline model, which excludes both the HTRP and FAF modules, achieves an OA of 55.39%, a Kappa coefficient of 0.5097, a SSIM of 0.805, and an inference time of 1131.0 ms. These results indicate that, without the integration of specialized modules, the model demonstrates limited capability in both classification accuracy and structural preservation, particularly in the presence of challenging noisy conditions. Furthermore, the baseline configuration incurs a relatively high computational cost during inference.
The incorporation of the HTRP module in isolation yielded substantial and multifaceted improvements across all evaluation metrics when contrasted with the baseline configuration. As delineated in Table III, the OA experienced an increase from the baseline of 55.39% to 57.08%. Similarly, the Kappa coefficient, a measure of agreement adjusted for chance, rose from 0.5097 to 0.5283. The SSIM, indicative of perceived image quality and structural preservation, also showed enhancement, improving from 0.805 to 0.833. This finding strongly suggests that the HTRP module not only augments the model’s capacity to effectively leverage domain-specific relational knowledge, thereby fostering the extraction of more salient temporal and relational features crucial for accurate classification, but also introduces a significant enhancement in computational performance. The pronounced reduction in inference time can be primarily attributed to the hierarchical design inherent to the HTRP module. Such an architecture is strategically conceived to streamline the propagation of information throughout the model and to minimize redundant computational steps, likely by creating more efficient pathways for feature processing or by reducing the overall parameter load associated with these operations. This gain in efficiency is of particular importance for applications deployed in real-time environments or those constrained by limited computational resources, where rapid inference and decision-making are paramount. It should be noted, however, that although the HTRP module leads to a notable increase in classification accuracy, it achieves a slightly lower SSIM compared to the FAF module. This phenomenon may be attributed to the introduction of textual information, which, while advantageous for classification, can introduce subtle conflicts with the objective of maintaining signal structure. This observation underscores an inherent trade-off between the extraction of discriminative features and optimal denoising performance.
Independent integration of the FAF module produces an OA of 57.25%, a Kappa coefficient of 0.5306, and an SSIM of 0.857, the highest SSIM among all configurations tested. The inference time in this case is 1149.4 ms, which represents a slight increase relative to the baseline. This modest rise in computational cost is offset by the substantial gain in structural preservation and denoising quality. The FAF module’s frequency-aware fusion mechanism enables the model to capture both global and local features by effectively integrating high- and low-frequency components, thus ensuring that essential structural information is maintained even in the presence of noise. The trade-off between the minor increase in inference time and the significant improvement in SSIM underscores the practical value of the FAF module for scenarios where denoising quality is paramount.
When both HTRP and FAF modules are jointly deployed, the model achieves the best overall performance, with an OA of 58.10%, a Kappa coefficient of 0.5391, and an SSIM of 0.838. The inference time in this configuration is 783.8 ms, which remains markedly lower than the baseline and only marginally higher than that of the HTRP-only model. This outcome demonstrates the complementary strengths of the two modules: HTRP not only boosts discriminative feature extraction and classification accuracy, but also streamlines the inference process, while FAF contributes to superior signal reconstruction at a negligible computational cost. Importantly, the combined model delivers consistent improvements across all evaluation metrics while maintaining efficient inference, highlighting its suitability for deployment in both accuracy- and latency-sensitive environments.
In summary, the ablation studies provide comprehensive evidence regarding the individual and joint contributions of the HTRP and FAF modules. The results affirm that each module addresses distinct yet complementary aspects of the overall task. The HTRP module stands out for its dual benefits of performance enhancement and significant reduction in inference time, making it highly valuable for practical applications. The FAF module, while incurring a slight increase in computational demand, delivers the highest gains in denoising quality. Their integration achieves an optimal balance between classification accuracy, structural preservation, and computational efficiency, thereby ensuring the robustness and applicability of the model in real-world, noisy environments where both performance and efficiency are critical.
IV-J Parameter Sensitivity Analysis
To assess the impact of key parameters on our framework’s performance, we conducted a sensitivity analysis focusing on Top-K ($K$) selection, Decoder usage, and LLM choice. The experiments were performed on the RML16A dataset under a consistent 100-shot learning setting to ensure fair comparisons. The results and observations for each parameter are detailed below.
IV-J1 Top-K
While increasing $K$ can improve key metrics such as OA and Kappa by capturing richer and more diverse information, it also leads to an increase in computational load, resulting in longer inference times. The Top-K selection mechanism was introduced to mitigate token redundancy by selecting the most informative tokens, thereby enhancing model efficiency without sacrificing critical information. However, as $K$ increases, although more informative tokens are incorporated, the computational overhead during both training and inference correspondingly rises.
This study systematically analyzes the sensitivity of OA, Kappa, and inference time (seconds per batch) to different values of $K$, highlighting the inherent trade-off between classification performance and computational efficiency. As depicted in Fig. 7, the results demonstrate a clear trend: initially, increasing $K$ leads to notable improvements in both OA and Kappa due to the inclusion of richer contextual information. However, the inference time also increases steadily with larger $K$ values.
Notably, $K=7$ emerges as the optimal configuration, achieving the highest OA and Kappa while maintaining a reasonable inference time. Beyond this point, the marginal gains in performance diminish relative to the additional computational cost, indicating a point of diminishing returns. Therefore, careful selection of $K$ is crucial to balance the benefits of richer token representation against the practical constraints of computational resources.
Figure 7: Performance Comparison Across Different $K$ Values.
IV-J2 Decoder Usage
To further investigate the impact of the decoder module on the overall model performance, we conducted an ablation study in which the Transformer decoder was replaced with a Linear decoder. Experimental results indicate that substituting the Transformer decoder with a Linear decoder results in a slight yet consistent degradation in performance. Specifically, the OA decreased from 58.10% to 57.41%, corresponding to a relative decline of 0.69%. Similarly, the Kappa coefficient dropped from 0.5391 to 0.5220, representing a reduction of 0.0171.
These results underscore the pivotal role of the Transformer decoder in enhancing both classification accuracy and prediction consistency. Benefiting from its global modeling capability, the Transformer decoder effectively refines contextual representations and improves information reconstruction, thereby contributing to superior task performance. In contrast, the Linear decoder lacks such global modeling capacity, leading to suboptimal results. Therefore, retaining the Transformer decoder is essential for achieving optimal performance within the proposed classification framework.
IV-J3 LLM Choice
The choice of LLM directly influences the performance and efficiency of the proposed RadioLLM framework. To systematically evaluate the impact, we evaluated three representative LLMs: BERT [42], GPT-2 [43], and LLaMA3 [44] based on OA, Kappa coefficient, and inference time (measured in seconds per batch). The results, summarized in Table IV, highlight the trade-offs between classification effectiveness and computational efficiency.
TABLE IV: Performance Comparison Across Different LLMs.
LLM OA (%) Kappa Inference Time (s)
BERT 57.53 0.5332 1.7519
GPT-2 58.10 0.5391 0.9069
LLaMA3 58.67 0.5480 0.9319
As shown in Table IV, LLaMA3 achieves the highest OA and Kappa coefficient among the evaluated models, indicating superior classification performance. This can be attributed to its advanced architectural design, including optimized attention mechanisms and improved parameter efficiency, as well as its training on more recent and diverse datasets. These factors collectively enhance LLaMA3’s ability to model complex patterns in radio signal data effectively.
GPT-2 demonstrates a balanced trade-off between performance and efficiency. Although its OA and Kappa are slightly lower than those of LLaMA3, it achieves the fastest inference time, making it a practical choice for latency-sensitive applications.
In contrast, BERT exhibits the lowest classification performance and the highest inference time. The increased computational cost associated with BERT can be largely attributed to its bidirectional attention mechanism, which requires attending to both past and future tokens simultaneously during each layer’s computation. While this bidirectional context modeling benefits certain natural language understanding tasks, it introduces additional overhead that is less favorable for the lightweight inference requirements of radio signal classification.
Overall, these findings underscore the importance of selecting an appropriate LLM based on the specific constraints and priorities of the application, balancing accuracy and computational efficiency accordingly.
V Conclusion